# Three dimensional end-to-end simulation for kilonova emission from a black-hole neutron-star merger

### v1

@@ Abstract
We study long-term evolution of the matter ejected in a black-hole neutron-star (BH-NS) merger employing the results of a long-term numerical-relativity simulation and nucleosynthesis calculation, in which both dynamical and post-merger ejecta formation are consistently followed. In particular, we employ the results for the merger of a $1.35\,M_{\odot}$ NS and a $5.4\,M_{\odot}$ BH with the dimensionless spin of 0.75. We confirm the finding in the previous studies that thermal pressure induced by radioactive heating in the ejecta significantly modifies the morphology of the ejecta. We then compute the kilonova (KN) light curves employing the ejecta profile obtained by the long-term evolution. We find that our present BH-NS model results in a KN light curve that is fainter yet more enduring than that observed in AT2017gfo. This is due to the fact that the emission is primarily powered by the lanthanide-rich dynamical ejecta, in which a long photon diffusion time scale is realized by the large mass and high opacity. While the peak brightness of the KN emission in both the optical and near-infrared bands is fainter than or comparable to those of binary NS models, the time-scale maintaining the peak brightness is much longer in the near-infrared band for the BH-NS KN model. Our result indicates that a BH-NS merger with massive ejecta can observationally be identified by the bright and long lasting ($>$two weeks) near-infrared emission.
@@ 1 Introduction
Neutron star (NS) mergers are known to be among the most promising targets of the ground-based gravitational-wave (GW) detectors (LIGO: Aasi et al. 2015, Virgo: Acernese et al. 2015, KAGRA: Kuroda 2010) as well as one of the most important sources of high-energy astrophysical transients, such as gamma-ray bursts (GRB, Paczynski, 1991; Nakar, 2007; Berger, 2014; Abbott et al., 2017c), kilonovae (KN, Li & Paczynski, 1998; Kulkarni, 2005; Metzger et al., 2010; Kasen et al., 2013; Tanaka & Hotokezaka, 2013), jet heated cocoons (Nakar & Piran, 2017; Hamidani & Ioka, 2023a, b), and synchrotron flares (Nakar & Piran, 2011; Hotokezaka & Piran, 2015; Hotokezaka et al., 2018; Margalit & Piran, 2020). NS mergers are also considered to be important production sites of elements heavier than iron in the universe (Lattimer & Schramm, 1974; Eichler et al., 1989; Freiburghaus et al., 1999; Cowan et al., 2021). The first detection of GWs from a binary neutron star (BNS) merger (GW170817; Abbott et al. 2017a) and its multi-wavelength electromagnetic (EM) counterparts  (Abbott et al., 2017b) demonstrated that those simultaneous observations will provide a valuable opportunity to extend our knowledge of fundamental physics in the extreme (strongly self-gravitating, high-density, and high-temperature) environments.
@@
To extract the physical information from the observation of EM counterparts, accurate modeling of the light curves and spectra consistent with the source properties are crucial. Since the detection of GW170817, light curve modeling of EM counterparts, particularly, for KNe has been significantly developed in this decade. In particular, the studies by employing numerical-simulation-based/motivated ejecta profiles and by performing radiative transfer (RT) simulations with realistic heating rates and/or detailed opacity tables enable us to directly connect the properties of the progenitor binary to the observables (e.g., Kasen et al., 2013, 2015; Barnes et al., 2016; Wollaeger et al., 2018; Tanaka et al., 2018; Wu et al., 2019; Kawaguchi et al., 2018; Hotokezaka & Nakar, 2020; Kawaguchi et al., 2020; Korobkin et al., 2021; Bulla et al., 2021; Zhu et al., 2021; Barnes et al., 2021; Nativi et al., 2020; Kawaguchi et al., 2021; Wu et al., 2022; Just et al., 2022, 2023). Previous studies showed that the complex ejecta profile in the presence of the multiple ejecta components of different mass ejection processes induces significant spatial dependences in radioactive heating as well as strong geometrical effects in radiative transfer, which have great impacts on the resulting light curves (Kasen et al., 2015; Kawaguchi et al., 2018, 2020; Bulla, 2019; Zhu et al., 2020; Darbha & Kasen, 2020; Korobkin et al., 2021; Almualla et al., 2021; Kedia et al., 2023). Hence, the employment of the realistic ejecta profile consistently taking multiple ejecta components into account is essential for the accurate prediction of KN light curves.
@@
One of the important missing links for the accurate prediction of KNe is the long-term hydrodynamics evolution of ejecta after the formation. While the ejecta formation takes place on a time scale of $\lesssim 1$–10 s after the onset of a merger (Hayashi et al., 2022, 2023), the KN emission peaks in a much longer time scale of 0.1–10 d (Li & Paczynski, 1998; Kulkarni, 2005; Metzger et al., 2010; Kasen et al., 2013; Tanaka & Hotokezaka, 2013), at which the homologous expansion of ejecta has been achieved. Since ejected matter can be accelerated by the ejecta pressure gradient and interact with different ejecta components during these epochs, the ejecta profile at the time of KN emission is non-trivial just from the ejecta properties at the time of formation. In fact, Rosswog et al. (2014) and Grossman et al. (2014) performed pseudo-Newtonian hydrodynamics simulations for BNS mergers, and studied the long-term evolution of the dynamical ejecta component until it reached the homologously expanding phase. They found that the thermal pressure induced by radioactive heating in ejecta significantly changes the ejecta morphology (see also Foucart et al. (2021)).  Fernández et al. (2015) and Fernández et al. (2017) performed long-term simulations for BH-NS mergers to investigate the effect of the interplay between the dynamical and post-merger components and found that the interaction of the multiple ejecta components can modify the ejecta profile. Thus, to accurately predict KN light curves, it is also important to follow the hydrodynamics evolution of the multiple ejecta components until the homologously expanding phase.
@@
Recently, the development of numerical simulation techniques and the significant increase in the computational resources have enabled us to consistently follow the NS mergers from the onset of the merger to the time that ejecta formation saturates  (Kiuchi et al., 2022; Fujibayashi et al., 2023, 2020b; Shibata et al., 2021; Hayashi et al., 2022; Kiuchi et al., 2022; Fujibayashi et al., 2023; Hayashi et al., 2023; Kiuchi et al., 2023; Just et al., 2023; Gottlieb et al., 2023; Kiuchi et al., 2024). In this paper, we study the KN emission associated with a BH-NS merger employing the results obtained by the numerical-relativity (NR) simulation and nucleosynthesis calculation consistently following the entire ejecta formation from the merger (Hayashi et al., 2022, 2023; Wanajo et al., 2022). In particular, we focus on the KN emission from $\approx 1\,{\rm d}$ after the onset of the merger for the model of a large amount of dynamical ejecta with $\approx 0.04M_{\odot}$ in this paper. [^1] During the submission process of this paper, LIGO and Virgo have detected GWs plausibly from a BH-NS merger (The LIGO Scientific Collaboration et al., 2024b). The alert shows that the system is likely to contain a NS, and the mass of the other object is likely to be in between $3\,M_{\odot}$ and $5\,M_{\odot}$ with a 50% probability. The probability for the matter outside the remnant object to be present after the merger is also high ($>99\%$). Hence, the system can be by chance similar to the BH-NS model studied in this paper (a binary of a $1.35\,M_{\odot}$ NS with the radius of $\approx 13.2\,{\rm km}$ and a $5.4\,M_{\odot}$ BH with the dimensionless spin of 0.75). We note that the amount of the dynamical ejecta is broadly the same ($\approx 0.04\,M_{\odot}$) also for a BH-NS merger with the same NS mass, NS radius, and dimensionless BH spin but with a larger BH mass ($8.1\,M_{\odot}$(Hayashi et al., 2022). The result of this paper indicates that, if it is a BH-NS merger with significant amount of the dynamical ejecta formation, this GW event may be associated with a kilonova of which near-infrared emission is bright and long-lasting.
@@ 2 Method
The simulation code for the HD simulation is a 3D extension of the code developed in our previous studies (Kawaguchi et al., 2021, 2022, 2023). This code solves the relativistic Euler equations under a spherical coordinate system. In order to incorporate the effect of gravity, a fixed background metric for a non-rotating black hole expressed in isotropic coordinates is used. See Appendix A for the formulation of the basic equations. The effect of radioactive heating is incorporated in the same way as in the previous studies (Kawaguchi et al., 2021, 2022, 2023). See Appendix B for the method of particle tracing used to employ the nucleosynthesis results in the HD simulation. We note that the equatorial symmetry is imposed for the HD simulation following the setup of the NR simulation.
@@
For the equation of state (EOS), we consider both contributions from gas and radiation: the total pressure $P$ is given by $P=P_{\rm gas}+P_{\rm rad}$ with $P_{\rm gas}=n_{\rm B}k_{\rm B}T$ and $P_{\rm rad}=a_{\rm rad}T^{4}/3$, where $n_{\rm B}$, $T$, $k_{\rm B}$, and $a_{\rm rad}$ are the baryon number density, temperature, Boltzmann constant, and radiation density constant, respectively. Here, we simplified the gas pressure assuming that atoms are fully ionized with an electron fraction of unity, and the gas pressure is dominated by the contribution from electrons (since the average atomic mass number is expected to be much larger than unity). We note that, although this simplification may overestimate the gas pressure component, the contribution of the gas pressure is found to be nevertheless subdominant. In fact, we confirm that the resulting ejecta profiles as well as the KN light curves are essentially unchanged even if we employ the ideal-gas EOS with the adiabatic index of $\Gamma=4/3$, which corresponds to the case that the radiation pressure dominates.
@@
Note that the magnetic field effects are not taken into account in our present HD simulations. As a consequence, and due to the coarse grid resolution in the polar region, the relativistic jet outflow launched in the NR simulation is not well resolved in the present HD simulations. The previous study suggests that the presence of the jet may affect the ejecta profile and hence the KN light curves near the jet axis (Nativi et al., 2020; Klion et al., 2021). Since resolving the propagation of the relativistic jet in long-term three-dimensional simulations requires high computational costs, we leave the investigation of the effect of the jet for a future work.
@@
We employ the same time origin for the HD simulations as in the NR simulations. The uniform grids with $N_{\theta}$ and $N_{\phi}$ grid points are prepared for the polar angle $\theta$ and the longitudinal angle $\phi$, respectively. For the radial direction, the following non-uniform grid structure is employed; for a given $j$-th radial grid point
@@
where $r_{\rm in}$ and $r_{\rm out}$ denote the inner and outer radii of the computational domain, respectively, and $N_{r}$ denotes the total number of the radial grid points. In the present work, we employ $(N_{r},N_{\theta},N_{\phi})=(1024,64,128)$, and $r_{\rm in}$ and $r_{\rm out}$ are initially set to be $3,000\,{\rm km}$ and $10^{3}\,r_{\rm in}$, respectively. We confirm that this grid resolution is sufficiently high enough for our purpose of the study by checking the results of the ejecta profile and KN light curves being semi-quantitatively unchanged for the HD simulation with $(N_{r},N_{\theta},N_{\phi})=(512,32,64)$ (less than 10% and 3% difference in the total bolometric luminosity at 1 d and 2 d, respectively).
@@
The so-called Courant–Friedrichs–Lewy (CFL) condition restricts the time steps in the HD simulation to ensure the numerical stability. For our setup, the time interval should be approximately less than the smallest value among $\Delta r_{\rm min}/c$, $r_{\rm in}\Delta\theta_{\rm min}/c$, and $r_{\rm in}{\rm sin}\theta_{\rm min}\Delta\phi_{\rm min}/c$ with $\theta_{\rm min}$, $\Delta r_{\rm min}$, $\Delta\theta_{\rm min}$, and $\Delta\phi_{\rm min}$ being the minimum cell center value of the $\theta$ coordinate and the minimum cell sizes of $r$, $\theta$, and $\phi$ directions, respectively. For the present grid setup, the most strict constraint comes from the last condition of $r_{\rm in}{\rm sin}\theta_{\rm min}\Delta\phi_{\rm min}/c$, and this restricts the time interval to be so small that the computational costs becomes practically quite high. To relax this condition, we average over the conservative variables of hydrodynamics in the direction of $\phi$ for all the cells located in $\theta\leq\theta_{\rm c}$ for each sub-step of the evolution. By this prescription, the HD simulation is kept numerically stable if the time interval is within $r_{\rm in}{\rm sin}\theta_{\rm c}\Delta\phi_{\rm min}/c$. For the present study, we choose $\theta_{\rm c}$ to be $\pi/24$, while we confirm that the resulting LCs are essentially unchanged even if we employ $\theta_{\rm c}=\pi/12$.
@@
The light curves of KNe are calculated using a wavelength-dependent RT simulation code (Tanaka & Hotokezaka, 2013; Tanaka et al., 2017, 2018; Kawaguchi et al., 2020, 2021). In this code, the photon transfer is simulated by a Monte Carlo method for given ejecta profiles composed of the density, velocity, and elemental abundance under the assumption of the homologous expansion. The time-dependent thermalization efficiency is taken into account following an analytic formula derived by Barnes et al. (2016). The ionization and excitation states are determined under the assumption of the local thermodynamic equilibrium (LTE) by using the Saha’s ionization and Boltzmann excitation equations.
@@
A three-dimensional cylindrical grid is applied for storing the local elemental abundances and radioactive heating rate as well as for solving the local temperature and opacity. The 50, 50, and 32 cells are set to the cylindrical radius, vertical, and longitudinal directions, which cover the domain with the coordinate ranges of $(0,0.6\,ct)$, $(0,0.6\,ct)$, and $(0,2\pi)$, respectively. We confirm that the resulting light curves are unchanged by changing each cell numbers from 50, 50, and 32 cells to $40$, $40$, and $28$ cells or changing the maximum cylindrical radius and vertical coordinate ranges from $0.6\,ct$ to $0.75\,ct$.

### v2

@@ Abstract
We study long-term evolution of the matter ejected in a black-hole neutron-star (BH-NS) merger employing the results of a long-term numerical-relativity simulation and nucleosynthesis calculation, in which both dynamical and post-merger ejecta formation is consistently followed. In particular, we employ the results for the merger of a $1.35\,M_{\odot}$ NS and a $5.4\,M_{\odot}$ BH with the dimensionless spin of 0.75. We confirm the finding in the previous studies that thermal pressure induced by radioactive heating in the ejecta significantly modifies the morphology of the ejecta. We then compute the kilonova (KN) light curves employing the ejecta profile obtained by the long-term evolution. We find that our present BH-NS model results in a KN light curve that is fainter yet more enduring than that observed in AT2017gfo. This is due to the fact that the emission is primarily powered by the lanthanide-rich dynamical ejecta, in which a long photon diffusion time scale is realized by the large mass and high opacity. While the peak brightness of the KN emission in both the optical and near-infrared bands is fainter than or comparable to those of binary NS models, the time-scale maintaining the peak brightness is much longer in the near-infrared band for the BH-NS KN model. Our result indicates that a BH-NS merger with massive ejecta can observationally be identified by the long lasting ($>$two weeks) near-infrared emission.
@@ 1 Introduction
Neutron star (NS) mergers are known to be among the most promising targets of the ground-based gravitational-wave (GW) detectors (LIGO: Aasi et al. 2015, Virgo: Acernese et al. 2015, KAGRA: Kuroda 2010) as well as one of the most important sources of high-energy astrophysical transients, such as gamma-ray bursts (GRB, Paczynski, 1991; Nakar, 2007; Berger, 2014; Abbott et al., 2017c), kilonovae (KNe, Li & Paczynski, 1998; Kulkarni, 2005; Metzger et al., 2010; Kasen et al., 2013; Tanaka & Hotokezaka, 2013), jet heated cocoons (Nakar & Piran, 2017; Hamidani & Ioka, 2023a, b), and synchrotron flares (Nakar & Piran, 2011; Hotokezaka & Piran, 2015; Hotokezaka et al., 2018; Margalit & Piran, 2020). NS mergers are also considered to be important production sites of elements heavier than iron in the universe (Lattimer & Schramm, 1974; Eichler et al., 1989; Freiburghaus et al., 1999; Cowan et al., 2021). The first detection of GWs from a binary neutron star (BNS) merger (GW170817; Abbott et al. 2017a) and its multi-wavelength electromagnetic (EM) counterparts  (Abbott et al., 2017b) demonstrated that those simultaneous observations will provide a valuable opportunity to extend our knowledge of fundamental physics in the extreme (strongly self-gravitating, high-density, and high-temperature) environments.
@@
To extract the physical information from the observation of EM counterparts, accurate modeling of the light curves and spectra consistent with the source properties are crucial. Since the detection of GW170817, light curve modeling of EM counterparts, particularly, for KNe has been significantly developed in this decade. In particular, the studies by employing numerical-simulation-based/motivated ejecta profiles and by performing radiative transfer (RT) simulations with realistic heating rates and/or detailed opacity tables enable us to directly connect the properties of the progenitor binary to the observables (e.g., Kasen et al., 2013, 2015; Barnes et al., 2016; Wollaeger et al., 2018; Tanaka et al., 2018; Wu et al., 2019; Kawaguchi et al., 2018; Hotokezaka & Nakar, 2020; Kawaguchi et al., 2020; Korobkin et al., 2021; Bulla et al., 2021; Zhu et al., 2021; Barnes et al., 2021; Nativi et al., 2020; Kawaguchi et al., 2021; Wu et al., 2022; Just et al., 2022; Just et al., 2023). Previous studies showed that the complex ejecta profile in the presence of the multiple ejecta components of different mass ejection processes induces significant spatial dependences in radioactive heating as well as strong geometrical effects in radiative transfer, which have great impacts on the resulting light curves (Kasen et al., 2015; Kawaguchi et al., 2018; Kawaguchi et al., 2020; Bulla, 2019; Zhu et al., 2020; Darbha & Kasen, 2020; Korobkin et al., 2021; Almualla et al., 2021; Kedia et al., 2023; Collins et al., 2023; Shingles et al., 2023; Collins et al., 2024). Hence, the employment of the realistic ejecta profile consistently taking multiple ejecta components into account is essential for the accurate prediction of KN light curves.
@@
One of the important missing links for the accurate prediction of KNe is the long-term hydrodynamics evolution of ejecta after the formation. While the ejecta formation takes place on a time scale of $\lesssim 1$–10 s after the onset of a merger (Hayashi et al., 2022, 2023), the KN emission peaks in a much longer time scale of 0.1–10 d (Li & Paczynski, 1998; Kulkarni, 2005; Metzger et al., 2010; Kasen et al., 2013; Tanaka & Hotokezaka, 2013), at which the homologous expansion of ejecta has been achieved. Since ejected matter can be accelerated by the pressure gradient in it and interact with different ejecta components during these epochs, the ejecta profile at the time of the KN emission is non-trivial just from the ejecta properties at the time of formation. In fact, Rosswog et al. (2014) and Grossman et al. (2014) performed pseudo-Newtonian hydrodynamics simulations for BNS mergers, and studied the long-term evolution of the dynamical ejecta component until they reached the homologously expanding phase. They found that the thermal pressure induced by radioactive heating in ejecta significantly changes the ejecta morphology (see also Foucart et al. 2021).  Fernández et al. (2015) and Fernández et al. (2017) performed long-term simulations for BH-NS mergers to investigate the effect of the interplay between the dynamical and post-merger components and found that the interaction of the multiple ejecta components can modify the ejecta profile. Thus, to accurately predict KN light curves, it is also important to follow the hydrodynamics evolution of the multiple ejecta components until the homologously expanding phase.
@@
Recently, the development of numerical simulation techniques and the significant increase in the computational resources have enabled us to consistently follow the NS mergers from the onset of the merger up to the time that ejecta formation saturates  (Kiuchi et al., 2022; Fujibayashi et al., 2023, 2020b; Shibata et al., 2021; Hayashi et al., 2022; Kiuchi et al., 2022; Fujibayashi et al., 2023; Hayashi et al., 2023; Kiuchi et al., 2023; Just et al., 2023; Gottlieb et al., 2023b; Kiuchi et al., 2024). In this paper, we study the KN emission associated with a BH-NS merger employing the results obtained by the numerical-relativity (NR) simulation and nucleosynthesis calculation consistently following the entire ejecta formation from the merger (Hayashi et al., 2022, 2023; Wanajo et al., 2022). In particular, we focus on the KN emission from $\approx 1\,{\rm d}$ after the onset of the merger for the model of a large amount of dynamical ejecta with $\approx 0.04M_{\odot}$ in this paper.
@@ 2 Method
The simulation code for the HD simulation is a 3D extension of the code developed in our previous studies (Kawaguchi et al., 2021, 2022, 2023). This code solves the relativistic Euler equations under a spherical coordinate system. In order to incorporate the effect of gravity, a fixed background metric for a non-rotating black hole expressed in isotropic coordinates is used (see Appendix A for the formulation of the basic equations). The effect of radioactive heating is incorporated by considering the neutrino loss [^1] The experimentally evaluated neutrino-energy loss for each $\beta$-decay is adopted from the ENDF/B-VIII.0 library (Brown et al., 2018). For those with no experimental evaluation (relevant to the entire phase of $r$-processing), the fraction of neutrino-energy loss is assumed to be 0.4 according to Hotokezaka et al. (2016). in the same way as in the previous studies (see Appendix A and  Kawaguchi et al. 2021, 2022, 2023; see also Appendix B for the method of particle tracing used to employ the nucleosynthesis results in the HD simulation). We note that the equatorial symmetry is imposed for the HD simulation following the setup of the NR simulation.
@@
For the equation of state (EOS), we consider both contributions from gas and radiation: the total pressure $P$ is given by $P=P_{\rm gas}+P_{\rm rad}$ with $P_{\rm gas}=n_{\rm B}k_{\rm B}T/\mu$ and $P_{\rm rad}=a_{\rm rad}T^{4}/3$, where $n_{\rm B}$, $T$, $k_{\rm B}$, $\mu$, and $a_{\rm rad}$ are the baryon number density, temperature, Boltzmann constant, mean molecular weight, and radiation density constant, respectively. Here, we simplified the gas pressure assuming that atoms are fully ionized with $\mu=1$, and the gas pressure is dominated by the contribution from electrons (since the average atomic mass number is expected to be much larger than unity). We note that, although this simplification may overestimate the gas pressure component, the contribution of the gas pressure is found to be nevertheless subdominant. In fact, we confirm that the resulting ejecta profiles as well as the KN light curves are essentially unchanged even if we employ the ideal-gas EOS with the adiabatic index of $\Gamma=4/3$, which corresponds to the case that the radiation pressure dominates.
@@
Note that the magnetic field effects are not taken into account in our present HD simulations. As a consequence, and due to the coarse grid resolution in the polar region, the collimated relativistic jet launched in the NR simulation is not well resolved in the present HD simulations. The previous study suggests that the presence of the jet may affect the ejecta profile and hence the KN light curves near the jet axis (Nativi et al., 2020; Klion et al., 2021). Since resolving the propagation of the collimated relativistic jet in long-term three-dimensional simulations requires high computational costs, we leave the investigation of the effect of the jet for future work.
@@
We employ the same time origin for the HD simulations as in the NR simulations. The uniform grid with $N_{\theta}$ and $N_{\phi}$ grid-cells along the polar angle $\theta$ and the longitudinal angle $\phi$, respectively, is prepared. For the radial direction, the following non-uniform grid structure is employed; for a given $j$-th radial grid-node
@@
where $r_{\rm in}$ and $r_{\rm out}$ denote the inner and outer radii of the computational domain, respectively, and $N_{r}$ denotes the total number of the grid-cells along the radial direction. In the present work, we employ $(N_{r},N_{\theta},N_{\phi})=(1024,64,128)$, and $r_{\rm in}$ and $r_{\rm out}$ are initially set to be $3,000\,{\rm km}$ and $10^{3}\,r_{\rm in}$, respectively. We confirm that this grid resolution is sufficiently high enough for our purpose of the study by checking the results of the ejecta profile and KN light curves being semi-quantitatively unchanged for the HD simulation with $(N_{r},N_{\theta},N_{\phi})=(512,32,64)$ (less than 10% and 3% difference in the total bolometric luminosity at 1 d and 2 d, respectively).
@@
The so-called Courant–Friedrichs–Lewy (CFL) condition restricts the time steps in the HD simulation to ensure the numerical stability. For our setup, the time interval should be approximately less than the smallest value among $\Delta r_{\rm min}/c$, $r_{\rm in}\Delta\theta_{\rm min}/c$, and $r_{\rm in}{\rm sin}\theta_{\rm min}\Delta\phi_{\rm min}/c$ with $\theta_{\rm min}$, $\Delta r_{\rm min}$, $\Delta\theta_{\rm min}$, and $\Delta\phi_{\rm min}$ being the minimum cell center value of the $\theta$ coordinate and the minimum cell sizes of $r$, $\theta$, and $\phi$ directions, respectively. For the present grid setup, the most strict constraint comes from the last condition of $r_{\rm in}{\rm sin}\theta_{\rm min}\Delta\phi_{\rm min}/c$, and this restricts the time interval to be so small that the computational costs becomes practically quite high. To relax this condition, we average over the conservative variables of hydrodynamics in the direction of $\phi$ for all the cells located in $\theta\leq\theta_{\rm c}$ for each sub-step of the evolution (see Hirai & Podsiadlowski (2022) for the similar prescription). By this prescription, the HD simulation is kept numerically stable if the time interval is within $r_{\rm in}{\rm sin}\theta_{\rm c}\Delta\phi_{\rm min}/c$. For the present study, we choose $\theta_{\rm c}$ to be $\pi/24$, while we confirm that the resulting LCs are essentially unchanged even if we employ $\theta_{\rm c}=\pi/12$.
@@
The light curves of KNe are calculated using a wavelength-dependent RT simulation code (Tanaka & Hotokezaka, 2013; Tanaka et al., 2017, 2018; Kawaguchi et al., 2020, 2021). In this code, the photon transfer is simulated by a Monte Carlo method for given ejecta profiles composed of the density, velocity, and elemental abundance under the assumption of the homologous expansion. The time-dependent thermalization efficiency is taken into account following an analytic formula derived by Barnes et al. (2016). In our RT code, the local gas temperature at each time step is calculated as $T_{\rm gas}=(u_{\rm rad}/a_{\rm rad})^{1/4}$, i.e., under the assumption of local thermodynamic equilibrium (LTE) using the Stefan-Boltzmann law with the radiation energy density, $u_{\mathrm{rad}}$, obtained by the RT simulation. Then, the ionization and excitation states are determined from the gas density and temperature by using the Saha’s ionization and Boltzmann excitation equations.
@@
A three-dimensional cylindrical grid is applied for storing the local elemental abundances and radioactive heating rate as well as for solving the temperature and opacity. The 50, 50, and 32 cells are set to the cylindrical radius, vertical, and longitudinal directions, which cover the domain with the coordinate ranges of $(0,0.6\,ct)$, $(0,0.6\,ct)$, and $(0,2\pi)$, respectively. We confirm that the resulting light curves are unchanged by changing each cell numbers from 50, 50, and 32 cells to $40$, $40$, and $28$ cells or changing the maximum cylindrical radius and vertical coordinate ranges from $0.6\,ct$ to $0.75\,ct$.